OpenStudy (anonymous):

Create a quadratic polynomial function f(x) and a linear binomial in the form (x − a). Part 1. Show all work using long division to divide your polynomial by the binomial. Part 2. Show all work to evaluate f(a) using the function you created. Part 3. Use complete sentences to explain how the remainder theorem is used to determine whether your linear binomial is a factor of your polynomial function

OpenStudy (anonymous):

my functions are (x^3 - 3x^2 - x + 3), and (x - 1).

OpenStudy (amistre64):

long division hasnt change since the 3rd grade ....

OpenStudy (anonymous):

@dan815 @Danielp @marissalovescats

OpenStudy (anonymous):

but it is using polynomials, just a tad harder

OpenStudy (amistre64):

its exactly the same hardness as normal :/ -------------------- x - 1 | x^3 - 3x^2 - x + 3 how many times does x go into x^3?

OpenStudy (anonymous):

3 times :P

OpenStudy (amistre64):

x*x^2 = x^3 .... id say x^2 times

OpenStudy (anonymous):

x*x=x^2 then x^2*x=x^3 ohhhhhhhh xD

OpenStudy (amistre64):

x^2 -------------------- x - 1 | x^3 - 3x^2 - x + 3 now we mulitply, and subtract, and pull down whats left x^2 -------------------- x - 1 | x^3 - 3x^2 - x + 3 -(x^3 - x^2) ------------ -2x^2 - x + 3 and repeat ... how many times does x go into -2x^2?

OpenStudy (anonymous):

2 times?

OpenStudy (amistre64):

almost :) x*-2x = -2x^2, lets go with -2x times

OpenStudy (anonymous):

oooh i was debating whether -2 or 2...

OpenStudy (amistre64):

x^2 - 2x -------------------- x - 1 | x^3 - 3x^2 - x + 3 now we mulitply, and subtract, and pull down whats left x^2 - 2x -------------------- x - 1 | x^3 - 3x^2 -(x^3 - x^2) ------------ -2x^2 - x + 3 -(-2x^2+2x) ------------ -3x + 3 and repeat ... how many times does x go into -3x?

OpenStudy (anonymous):

-3 times

OpenStudy (amistre64):

yes, this time its just -3 :) -3*x = -3x soo what do we do?

OpenStudy (anonymous):

divide by -3?

OpenStudy (amistre64):

notice that we are simply playing with first terms such that: x * k = x^3 k = x^3/x = x^2 or again x * k = -2x^2 k = -2x^2/x = -2x

OpenStudy (amistre64):

and no we dont divide by -3 ... we use it in the same way weve always used it when doing longhand division .... the same way i demonstrated the two times before

OpenStudy (amistre64):

multply, and subtract ....

OpenStudy (anonymous):

@amistre64 so do x-1*-3? im confused...

OpenStudy (amistre64):

of course ...

OpenStudy (anonymous):

mmk little confused spo was checking

OpenStudy (anonymous):

so x*-3 is -3x and -1*-3=3

OpenStudy (amistre64):

refresh with numbers if its been too long :) 00 ------- 52 | 1245 124/52 = 2.xxx , use 2, 2(52) = 104 002 ------- 52 | 1245 -104 ---- 205 , drop down the 5 and repeat ...

OpenStudy (anonymous):

so 20*x-1

OpenStudy (anonymous):

20x-20

OpenStudy (amistre64):

not sure why youre going with 20 ... -3(x-1) = -3x+3 .... then subtract it from the rest

OpenStudy (anonymous):

oh you said drop the 5 so... oh

OpenStudy (amistre64):

forgot something so reedit to include that -3 again :) how many times does x go into -3x? -3 times soo x^2 - 2x - 3 -------------------- x - 1 | x^3 - 3x^2 - x + 3 now we mulitply, and subtract, and pull down whats left x^2 - 2x - 3 -------------------- x - 1 | x^3 - 3x^2 -(x^3 - x^2) ------------ -2x^2 - x + 3 -(-2x^2+2x) ------------ -3x + 3 -(-3x + 3) --------- 0, our remainder is simply zero now

OpenStudy (anonymous):

so what would we do for part 2?

OpenStudy (amistre64):

librarys closing so youre on your own ... or at least im on my own :) good luck

OpenStudy (anonymous):

Oh ok thx